OpenStudy (anonymous):

The graph below shows the solution to a system of inequalities: Solid line joining ordered pairs 0, 3.75 and 15, 0. Shade the portion of the graph above the line in the first and second quadrant Which of the following inequalities is modeled by the graph? 5x + 4y ≥ 15; y ≥ 0 5x - 4y ≥ 15; y ≥ 0 5x + 4y ≤ 15; y ≥ 0 -5x - 4y ≥ 15; y ≥ 0

OpenStudy (anonymous):

@Elsa213

OpenStudy (anonymous):

@lacrosseplayer22

Elsa213 (elsa213):

1) find the slope of the line through these 2 points; 2) insert this slope and the coordinates of either one of the given points into the point-slope form of the equation of a line: y-y0 = m (x-x0). What part of this can you do yourself and what part requires some guidance? 4) Where did that come from? I reduced Nelson's -3.75/15 and obtained the simple proper fraction -1/4. So, DG, are we in agreement or no? Be certain that you're comfortable with the slope calculation as well as the result.So, your slope is m=-1/4. Please choose one of the points (which one you choose does not matter at all).

Elsa213 (elsa213):

I got this from: http://openstudy.com/study#/updates/52d6059ee4b0274b88c56e7a :)

OpenStudy (anonymous):

wait so what was the answer?

Elsa213 (elsa213):

Wut do u think the answer is?

OpenStudy (anonymous):

Uhh C?

Elsa213 (elsa213):

5x + 4y ≤ 15; y ≥ 0 :)

OpenStudy (anonymous):

Can you help me with a few more?

Elsa213 (elsa213):

im not sure. I not great in math. Probably @lacrosseplayer22 can help

OpenStudy (anonymous):

Ok, Ill just ask the next question :)

Elsa213 (elsa213):

ok :) Also welcome to openstudy :D Hope you enjoy

OpenStudy (anonymous):

Bess is trying to solve a puzzle where she has to figure out two numbers, x and y. The sum of twice x and 3 times y is less than 6. Also, y must be greater than 1 less than twice the x. Which graph represents the possible solutions? Graph the inequalities 2 x plus 3 y less than 6 and y greater than 2 x minus 1 Graph the inequalities 2 x minus 3 y less than 6 and y greater than 2 x minus 1 Graph the inequalities 3 x minus 2 y less than 6 and y greater than 3 x minus 2 Graph the inequalities 3 x minus 5 y less than 6 and y greater than 3 x plus 2

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

What was the answer the the Bess one??

OpenStudy (secret-ninja):

The answer to the Bess one is A. @shybearr15